OpenStudy (ajr262636):

the pair of points (6,y) and (10,-1) lie on a line with slope 1/4. what is the value of y?

OpenStudy (alexandervonhumboldt2):

look: \[\frac{ y+1 }{ 6-10 }=1/4\]

OpenStudy (alexandervonhumboldt2):

\[y+1=1/4*-4\] \[y+1=-1\] \[y=-2\] lets check: (-2+1)/(6-10)=(-1)/(-4)=1/4 so y=-2

OpenStudy (ajr262636):

can you help me with this question? What is the slope of a vertical line?